OpenStudy (anonymous):

Solve using substitution method: 9x - 7y = 8

OpenStudy (anonymous):

I don't know how to do substitution.

OpenStudy (anonymous):

there's supposed be another equation....

razor99 (razor99):

is that algebra.

razor99 (razor99):

yeah hey where is the pther equation.

OpenStudy (anonymous):

oh woops 9x - 7y = 8 18x - 14y = -7

Directrix (directrix):

9x - 7y = 8 9x = 8 + 7y x = (8 + 7y)/9 Substitute into the second equation: 18 [(8 + 7y)/9] - 14y = -7 2(8 + 7y) - 14 y = -7 16 + 14y - 14 y = -7 16 = -7 which cannot happen There is no solution to this system of equations. The graphs of the two lines are parallel and there is no common solution.

Directrix (directrix):

See graph at the link for visual confirmation. http://www.wolframalpha.com/input/?i=%289x+-+7y+%3D+8%29+and+%2818x+-+14y+%3D+-7%29